# LogogramNLP: Comparing Visual and Textual Representations of Ancient Logographic Writing Systems for NLP

### v1

@@ 1 Introduction
The challenges associated with NLP for ancient logographic languages mainly come from two aspects. First, for many ancient languages, most available data sources are in visual forms, consisting of untranscribed photographs or hand-drawn copies (i.e., lineart). Adopting the conventional NLP pipeline, which requires converting visual representations into symbolic text, is therefore not straightforward: automatic transcriptions are often noisy due to data scarcity, while manual transcriptions are labor-intensive and require domain expertise. Some logographic writing systems, such as Old Chinese, even include symbol inventories that remain not fully mapped to Unicode (depicted in Figure 1).
@@
In this paper, we attempt to answer the following questions: (1) Can we effectively apply NLP toolkits, such as classifiers, machine translation systems, and syntactic parsers, to visual representations of logographic writing systems? (2) Does this strategy allow for better transfer from pre-trained models and lead to better performance? Additionally, as shown in Figure 1, many logographic languages have multiple partially processed representations, including artifact photographs, hand-copied lineart, Unicode, Latin transliteration, and normalization—we also aim to empirically investigate the extent to which various representations at each stage, including textual and visual modalities, facilitate effective fine-tuning of downstream NLP systems.
@@ 2 Dataset: Languages, Tasks and Challenges
Our benchmark consists of four representative ancient languages—Linear A, Egyptian hieroglyphic, Cuneiform, and Bamboo script (§ 2.1). [^1] Bamboo scripts usually combine Seal scripts and Clerical scripts. Each language is associated with a unique writing system and unique challenges. We refer the readers to Appendix A for data collection and cleaning details. Our benchmark covers three tasks: machine translation, dependency parsing, and attribute classification (§ 2.2).
@@ 2.1 Logographic Languages
A major characteristic of logographic languages is that the size of symbol inventories is significantly larger than that in alphabetic languages such as Ancient Greek (24 letters) or Modern English (26 letters). A summary of different representations of the languages of our interest is shown in Figure 2, and Table 2 summarizes the current status of each language.
@@ 2.1.1 Visual Representations
Since ancient scripts did not consistently adhere to a left-to-right writing order, breaking down multi-line documents into images of single-line text is nontrivial. These historical data, therefore, need additional processing to be machine-readable. Figure 3 shows examples of different processing strategies. We summarize the approaches we used in building the dataset as follows:
@@ 2.1.2 Textual Representations
For Egyptian hieroglyphs, most characters are encoded in Unicode, but there is no standard encoding for “stacking” multiple glyphs vertically (Figure 3). Therefore, we do not include the Unicode text for our ancient Egyptian data as they are not available.
@@ 2.2 Tasks
Our benchmark covers three tasks (Table 1): translation, dependency parsing, and attribute classification. The model performance on these tasks reflects various aspects of ancient language understanding. To better understand the information loss when using a pipeline approach, we also report performance using this method: predicting the transliteration first and using the noisy predicted transliteration for downstream tasks.
@@ 3 Methods
In this section, we will describe feature encoding methods (§ 3.1) for both visual and textual inputs, as well as task-specific layers (§ 3.2) for each task we consider.
@@ 3.1 Feature Encoding
The majority of past work on cross-lingual transfer has focused on using Latin transliteration as the proxy to transfer knowledge from high-resource to low-resource languages Pires et al. (2019); Fang et al. (2020). Following this line of work, we input latinization representations to mBERT (Devlin et al., 2018) to obtain the embeddings of the ancient languages.
@@ 4 Experiments and Analysis
We describe our general model fine-tuning approach in § 4.1 and analyze model performance on the aforementioned tasks in the succeeding subsections.
@@
Textual BPE w/ ext vocab2 Unicode BERT n/a $0^{**}$ $0^{**}$ $0^{**}$ n/a 74.85
@@ 4.1 General Experimental Setup
We use Adam (Kingma and Ba, 2015) as the optimizer for all models, with an initial learning rate specified in Table 3. We use early stopping when the validation loss fails to improve for ten evaluation intervals (1000 iteration per interval). For data without a standard test set, we run a fixed number of training iterations and report the performance on the validation set after the last iteration. All experiments are conducted on an NVIDIA-RTX A6000 GPU, and the training time ranges from 2 minutes to 50 hours, depending on the nature of the task and the size of the datasets. Unless otherwise specified, all parameters, including those in pre-trained models, are trainable without freezing. We summarize other configurations in Table 3.
@@ 4.2 Machine Translation
We compare the performance of the models on machine translation, where we translate ancient Egyptian (EGY), Akkadian (AKK), and Old Chinese (ZHO) into English (Table 4(a)). We find that the PIXEL-MT model consistently achieves the best BLEU score across the three languages, outperforming the second-best method by a large margin.
@@
As shown in Figure 4, the low BLEU scores for ZHO-EN translation is a result of the translation model failing to capture the meaning of the input, instead focusing on repeated formatting queues: e.g., “Confucius said: Those who...” Indeed, given that the topical domain of the ZHO-EN translation data is philosophical writing, achieving an accurate translation would be challenging even with a much larger set of parallel translations. For AKK-EN, we found that the overall quality to be quite good, despite the fact that errors in translating named entities appear more often than in standard MT tasks. This case study suggests that translation performance could improve further if we training using a custom target language (English) vocabulary. We also show more generated examples from the PIXEL-MT model in Appendix C.
@@ 4.3 Attribute Classification
Table 4(b) summarizes the performance of attribute classification with different features and models. As expected, the image features can work fairly well for some of these attribute classification tasks as many of the relevant features are visual (e.g., for time and location); but, are not generally as effective as textual input representations. By comparing BERT with latinized input and CANINE on Unicode, we find that when both accurate latinization and Unicode representations are available, latinization is the most informative feature—with the exception of time period classification for Akkadian. This exception is aligned with our understanding of Akkadian, as different Cuneiform characters are used across different time periods. Thus, in this case, Unicode can provide more clues for determining the time period of a sample. Note that the label distribution is not balanced for most ancient language attribution tasks. For more details, refer to Chen et al. (2024).
@@ 4.4 Dependency Parsing
We compare the dependency parsing performance of models with visual and textual encoders (Table 5). [^10] We only conduct experiments on Akkadian since it is the only language with off-the-shelf dependency annotations. While all models achieve quite high parsing accuracy, we find that models with visual encoders perform the best on both investigated corpora (RIAO and MCONG). During training, models taking visual input generally converge faster than their textual counterparts, which is in line with prior work (Salesky et al., 2023) that uses visual features for machine translation.
@@ 5 Ablation Study on OCR and Image Quality
As mentioned earlier, the majority of data from ancient times remain in the form of photographs. We first closely examine two different visual input representations for the ZHO-EN translation task, handcopied figure and photograph (§ 5.1). Next, we examine OCR performance on ancient logographic languages to gain better understanding of this bottleneck for current NLP pipelines (§ 5.2).
@@
For the ZHO-EN translation data, we have access to both photographs of the bamboo slips and handcopied textline figures (see the Bamboo script example in Figure 3 for reference). As shown in Table 6, the quality of the visual features significantly influences the translation accuracy—translations derived from photographs yield a low BLEU score of 2.09, whereas handcopied figures, which typically provide clearer and more consistent visual data, result in a higher BLEU score of 5.45. This result suggests that for models that perform implicit OCR as part of the translation process, the clarity of the source material is paramount.
@@ 6 Related work
Logography typically denotes a writing system in which each glyph represents a semantic value rather than a phonetic one, however, all the languages studied in our paper have at least some phonetic component based on the rebus principle. This paper emphasizes ancient logographies that (i) possess extensive glyph inventories; (ii) have feature glyphs with ambiguous transliterations or functional uses; and (iii) are low-resource with much of data remaining in photo formats (Caplice et al., 1991; Allen, 2000; Woodard, 2004). Existing research on logographic languages has predominantly focused on those well-resourced and still in use, such as Modern Chinese (Zhang and Komachi, 2018; Si et al., 2023), or used data that has already been carefully transcribed into Latin or annotated with extra semantic information (Wiesenbach and Riezler, 2019; Gutherz et al., 2023; Jiang et al., 2024). Our paper aims to address the gap in resources (by proposing new data) and methodologies (by adapting visual-only approaches) for encoding and analyzing ancient logographic languages, leading to more comprehensive understanding of historical linguistic landscapes.
@@ C.1 AKK - EN
Akkadian translation samples from the PIXEL-MT models is shown in Figure 7. We showcase the first 10 examples from the validation set, we find that some annotation error present for example #2 and #6.
@@ C.2 EGY - EN
Ancient Egyptian translation samples from the PIXEL-MT models is shown in Figure 8.
@@ C.3 ZHO - EN
Old Chinese translation samples from the PIXEL-MT models is shown in Figure 9.

### v2

@@ 1 Introduction
The challenges associated with NLP for ancient logographic languages mainly come from two aspects. First, for many ancient languages, most available data sources are in visual forms, consisting of untranscribed photographs or hand-drawn copies (i.e., lineart). Adopting the conventional NLP pipeline, which requires converting visual representations into symbolic text, is therefore not straightforward: automatic transcriptions are often noisy due to data scarcity, while manual transcriptions are labor-intensive and require domain expertise. Some logographic writing systems, such as Old Chinese, even include symbol inventories that remain not fully mapped to Unicode (depicted in Figure˜1).
@@
In this paper, we attempt to answer the following questions: (1) Can we effectively apply NLP toolkits, such as classifiers, machine translation systems, and syntactic parsers, to visual representations of logographic writing systems? (2) Does this strategy allow for better transfer from pre-trained models and lead to better performance? Additionally, as shown in Figure˜1, many logographic languages have multiple partially processed representations, including artifact photographs, hand-copied lineart, Unicode, Latin transliteration, and normalization—we also aim to empirically investigate the extent to which various representations at each stage, including textual and visual modalities, facilitate effective fine-tuning of downstream NLP systems.
@@ 2 Dataset: Languages, Tasks and Challenges
Our benchmark consists of four representative ancient languages—Linear A, Egyptian hieroglyphic, Cuneiform, and Bamboo script (§˜2.1). [^1] Bamboo scripts usually combine Seal scripts and Clerical scripts. Each language is associated with a unique writing system and unique challenges. We refer the readers to Appendix˜A for data collection and cleaning details. Our benchmark covers three tasks: machine translation, dependency parsing, and attribute classification (§˜2.2).
@@ 2.1 Logographic Languages
A major characteristic of logographic languages is that the size of symbol inventories is significantly larger than that in alphabetic languages such as Ancient Greek (24 letters) or Modern English (26 letters). A summary of different representations of the languages of our interest is shown in Figure˜2, and Table˜2 summarizes the current status of each language.
@@ 2.1.1 Visual Representations
Since ancient scripts did not consistently adhere to a left-to-right writing order, breaking down multi-line documents into images of single-line text is nontrivial. These historical data, therefore, need additional processing to be machine-readable. Figure˜3 shows examples of different processing strategies. We summarize the approaches we used in building the dataset as follows:
@@ 2.1.2 Textual Representations
For Egyptian hieroglyphs, most characters are encoded in Unicode, but there is no standard encoding for “stacking” multiple glyphs vertically (Figure˜3). Therefore, we do not include the Unicode text for our ancient Egyptian data as they are not available.
@@ 2.2 Tasks
Our benchmark covers three tasks (Table˜1): translation, dependency parsing, and attribute classification. The model performance on these tasks reflects various aspects of ancient language understanding. To better understand the information loss when using a pipeline approach, we also report performance using this method: predicting the transliteration first and using the noisy predicted transliteration for downstream tasks.
@@ 3 Methods
In this section, we will describe feature encoding methods (§˜3.1) for both visual and textual inputs, as well as task-specific layers (§˜3.2) for each task we consider.
@@ 3.1 Feature Encoding
The majority of past work on cross-lingual transfer has focused on using Latin transliteration as the proxy to transfer knowledge from high-resource to low-resource languages Pires et al. (2019); Amrhein and Sennrich (2020). Following this line of work, we input latinization representations to mBERT (Devlin et al., 2018) to obtain the embeddings of the ancient languages.
@@ 4 Experiments and Analysis
We describe our general model fine-tuning approach in §˜4.1 and analyze model performance on the aforementioned tasks in the succeeding subsections.
@@
Textual BPE w/ ext vocab2 Unicode BERT n/a $^{**} 00^{**}$ $0^{**}$ n/a 74.85
@@ 4.1 General Experimental Setup
We use Adam (Kingma and Ba, 2015) as the optimizer for all models, with an initial learning rate specified in Table˜3. We use early stopping when the validation loss fails to improve for ten evaluation intervals (1000 iteration per interval). For data without a standard test set, we run a fixed number of training iterations and report the performance on the validation set after the last iteration. All experiments are conducted on an NVIDIA-RTX A6000 GPU, and the training time ranges from 2 minutes to 50 hours, depending on the nature of the task and the size of the datasets. Unless otherwise specified, all parameters, including those in pre-trained models, are trainable without freezing. We summarize other configurations in Table˜3.
@@ 4.2 Machine Translation
We compare the performance of the models on machine translation, where we translate ancient Egyptian (EGY), Akkadian (AKK), and Old Chinese (ZHO) into English (Table˜4(a)). We find that the PIXEL-MT model consistently achieves the best BLEU score across the three languages, outperforming the second-best method by a large margin.
@@
As shown in Figure˜4, the low BLEU scores for ZHO-EN translation is a result of the translation model failing to capture the meaning of the input, instead focusing on repeated formatting queues: e.g., “Confucius said: Those who...” Indeed, given that the topical domain of the ZHO-EN translation data is philosophical writing, achieving an accurate translation would be challenging even with a much larger set of parallel translations. For AKK-EN, we found that the overall quality to be quite good, despite the fact that errors in translating named entities appear more often than in standard MT tasks. This case study suggests that translation performance could improve further if we training using a custom target language (English) vocabulary. We also show more generated examples from the PIXEL-MT model in Appendix˜C.
@@ 4.3 Attribute Classification
Table˜4(b) summarizes the performance of attribute classification with different features and models. As expected, the image features can work fairly well for some of these attribute classification tasks as many of the relevant features are visual (e.g., for time and location); but, are not generally as effective as textual input representations. By comparing BERT with latinized input and CANINE on Unicode, we find that when both accurate latinization and Unicode representations are available, latinization is the most informative feature—with the exception of time period classification for Akkadian. This exception is aligned with our understanding of Akkadian, as different Cuneiform characters are used across different time periods. Thus, in this case, Unicode can provide more clues for determining the time period of a sample. Note that the label distribution is not balanced for most ancient language attribution tasks. For more details, refer to Chen et al. (2024).
@@ 4.4 Dependency Parsing
We compare the dependency parsing performance of models with visual and textual encoders (Table˜5). [^10] We only conduct experiments on Akkadian since it is the only language with off-the-shelf dependency annotations. While all models achieve quite high parsing accuracy, we find that models with visual encoders perform the best on both investigated corpora (RIAO and MCONG). During training, models taking visual input generally converge faster than their textual counterparts, which is in line with prior work (Salesky et al., 2023) that uses visual features for machine translation.
@@ 5 Ablation Study on OCR and Image Quality
As mentioned earlier, the majority of data from ancient times remain in the form of photographs. We first closely examine two different visual input representations for the ZHO-EN translation task, handcopied figure and photograph (§˜5.1). Next, we examine OCR performance on ancient logographic languages to gain better understanding of this bottleneck for current NLP pipelines (§˜5.2).
@@
For the ZHO-EN translation data, we have access to both photographs of the bamboo slips and handcopied textline figures (see the Bamboo script example in Figure˜3 for reference). As shown in Table˜6, the quality of the visual features significantly influences the translation accuracy—translations derived from photographs yield a low BLEU score of 2.09, whereas handcopied figures, which typically provide clearer and more consistent visual data, result in a higher BLEU score of 5.45. This result suggests that for models that perform implicit OCR as part of the translation process, the clarity of the source material is paramount.
@@ 6 Related work
Logography typically denotes a writing system in which each glyph represents a semantic value rather than a phonetic one, however, all the languages studied in our paper have at least some phonetic component based on the rebus principle. This paper emphasizes ancient logographies that (i) possess extensive glyph inventories; (ii) have feature glyphs with ambiguous transliterations or functional uses; and (iii) are low-resource with much of data remaining in photo formats (Caplice and others, 1991; Allen, 2000; Woodard, 2004). Existing research on logographic languages has predominantly focused on those well-resourced and still in use, such as Modern Chinese (Zhang and Komachi, 2018; Si et al., 2023), or used data that has already been carefully transcribed into Latin or annotated with extra semantic information (Wiesenbach and Riezler, 2019; Gutherz et al., 2023; Jiang et al., 2024). Our paper aims to address the gap in resources (by proposing new data) and methodologies (by adapting visual-only approaches) for encoding and analyzing ancient logographic languages, leading to more comprehensive understanding of historical linguistic landscapes.
@@ C.1 AKK - EN
Akkadian translation samples from the PIXEL-MT models is shown in Figure˜7. We showcase the first 10 examples from the validation set, we find that some annotation error present for example #2 and #6.
@@ C.2 EGY - EN
Ancient Egyptian translation samples from the PIXEL-MT models is shown in Figure˜8.
@@ C.3 ZHO - EN
Old Chinese translation samples from the PIXEL-MT models is shown in Figure˜9.